I already have a 50% mixture. (It contains 7 scoops of A, - and 7 scoops of B.)
I want to take one scoop of the 50% mixture and and add enough B to it so that it will become 75%B and 25%A.
How do I proceed?

Answer by stanbon(75887)

I already have a 50% mixture. (It contains 7 scoops of A, - and 7 scoops of B.)
I want to take one scoop of the 50% mixture and and add enough B to it so that it will become 75%B and 25%A.
------------------------------
Equation:
Bstuff + Bstuff = Bstuff
-----------------------------
0.50*1scoop + 1*Bscoops = 0.75(1+B)
----
50*1 + 100B = 75*1 + 75B
25B = 25*1
B = 1 scoop (Amt of B to be added)
---------------------------
Cheers,
Stan H.
===========